OpenStudy (anonymous):

1/3 log x^2 + 1/3 log 27x=log 9

OpenStudy (anonymous):

Please show me the steps. I got 3 and I think thats incorrect.

OpenStudy (lgbasallote):

3 seems to be right :)

OpenStudy (anonymous):

@lgbasallote. Okay I checked my answer in wolfram and it says the answer should be -2 log^2(3)/log (9)-1. ??

OpenStudy (zepp):

Are these your steps?

OpenStudy (anonymous):

No. I got 3 and the above is another answer that was online. I need to know which one is correct and if its the other one I need to know how they got that.

OpenStudy (zepp):

I got 1/3. :|

OpenStudy (anonymous):

ok let me try it again...

OpenStudy (zepp):

http://puu.sh/txcq

OpenStudy (lgbasallote):

@zepp log(2) + log(x)?

OpenStudy (zepp):

log(27) + log(x)

OpenStudy (lgbasallote):

@zepp where did the 1/3 go again?

OpenStudy (zepp):

What 1/3? D:

OpenStudy (lgbasallote):

you know 1/3 log x + 1/3 log (27x) those 1/3

OpenStudy (lgbasallote):

im pretty sure you multiply them somewhere?

OpenStudy (zepp):

yeah

OpenStudy (zepp):

uh

OpenStudy (zepp):

http://puu.sh/txh1 that?

OpenStudy (lgbasallote):

\[\Large \log x^{\frac{2}{3}} + \log (3 x^{\frac{1}{3}}) = \log 9\] \[\Large \log (3 x^{\frac{2}{3}} + \frac{1}{3}) = \log 9\] \[\Large \log (3x) = \log 9\] \[3x = 9\] \[x = 3\]

OpenStudy (lgbasallote):

the 4th to the last line should be \[\Large \log (3x^{(\frac{2}{3} + \frac{1}{3})})\]

OpenStudy (accessdenied):

\[ \begin{align} \frac{1}{3} \log x^2 + \frac{1}{3} \log 27x &= \log 9 \\ \left( \frac{1}{3} \right) \left( \log x^2 + \log 27x \right) &= \log 9 \\ \log x^2 + \log 27x &= 3\log 9 \\ \log \:(x^2 \times 27x) &= 3\log 9 \\ \log 27x^3 &= 3\log 9 \\ 3\log 3x &= 3\log 9 \\ \log 3x &= \log 9 \\ 3x &= 9 \\ x &= 3 \end{align} \]

OpenStudy (lgbasallote):

how come align& doesnt work for me -_-

OpenStudy (accessdenied):

begin{align} \frac{1}{3} \log x^2 + \frac{1}{3} \log 27x &= \log 9 \\ \left( \frac{1}{3} \right) \left( \log x^2 + \log 27x \right) &= \log 9 \\ \log x^2 + \log 27x &= 3\log 9 \\ \log x^2 \times 27x &= 3\log 9 \\ \log 27x^3 &= 3\log 9 \\ 3\log 3x &= 3\log 9 \\ \log 3x &= \log 9 \\ 3x &= 9 \\ x &= 3 end{align} This is exactly how it was written inside the brackets, omitting the \ in front of begin and end to avoid magic

OpenStudy (accessdenied):

noticing & is in every line and a line break "\\" ends each line

OpenStudy (lgbasallote):

uhmm what does the & do? all the & aligns?

OpenStudy (accessdenied):

Yeah. In this case, I made it all line up on the equals sign

OpenStudy (lgbasallote):

ohh nice trick